OpenStudy (anonymous):

A. x = 2 B. x = 2.5 C. x = -1 D. x = 0.5

OpenStudy (anonymous):

So what's the question?

OpenStudy (anonymous):

x^2-3x=-5/4 Which choices are solutions to the following equation?

OpenStudy (anonymous):

Have you tried using the quadratic equation?

OpenStudy (anonymous):

no how ?

OpenStudy (anonymous):

this is the quadratic:\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

is it -1

OpenStudy (anonymous):

sorry it didnt come out right http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

no its going to be two values

OpenStudy (anonymous):

so you first want to set the given equation =0 kk? How do you do that?

OpenStudy (anonymous):

yea -1 and 2

OpenStudy (anonymous):

do wht .

OpenStudy (anonymous):

\[x ^{2}-3x+5/4\]

OpenStudy (anonymous):

all that =0

OpenStudy (anonymous):

wht ? confused .

OpenStudy (anonymous):

whts tht equation for .

OpenStudy (anonymous):

kk wait

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[ax ^{2} + bx + a =0 \] where you plug is a b c in the quadratic equation \[(-b \pm \sqrt{b ^{2}-4ac})/(2a)\]

OpenStudy (anonymous):

to get to values

OpenStudy (anonymous):

can you tell me the answer please .

OpenStudy (anonymous):

in this case a=1 b=-3 a=5/4

OpenStudy (anonymous):

umm just plug it in I gave you want to plug in

OpenStudy (anonymous):

it didnt give me no answer

OpenStudy (anonymous):

just plug in x =2, x = 2.5, x = -1 and x = 0.5 in the equation and see which value solves it. That will be your answer.

OpenStudy (anonymous):

im still aint geting no answer something wrong

OpenStudy (anonymous):

put x = 2 in \[x ^{2}-3x = -5/4\]

OpenStudy (anonymous):

you will get \[2^{2}-3*2 = -5/4 => 4-6 = -5/4=>-2=-5/4\]

OpenStudy (anonymous):

idk can i have the answer please .

OpenStudy (anonymous):

-2 is obviously not equal to -5/4. so x = 2 is not the answer. similarly put x = 2.5, x = -1, x = 0.5 in the equation and see which one solve the equation

OpenStudy (anonymous):

No, i can't do your problem for you. I can only show you how to do your problem

OpenStudy (anonymous):

well , i dont know how to do this so . its only right if you do .

OpenStudy (anonymous):

I showed you how to do it. just follow the steps.

OpenStudy (anonymous):

rd can you tell me how to put it in the calculator .

OpenStudy (anonymous):

Do you you know how to square things?\[X^2 \] ?

OpenStudy (anonymous):

yess that = 1/2

OpenStudy (anonymous):

no \[x^2 = x \times x\]

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

so just replace x with the given values and if the the equation makes sense then it works.

OpenStudy (anonymous):

Do you know how to multiply and subtract?

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

then just replace X with all the given values

OpenStudy (anonymous):

and if the equation makes sense then thats one of the answers

OpenStudy (anonymous):

so you sayinq 1/4 subtract all choices .

OpenStudy (anonymous):

how did you get 1/4

OpenStudy (anonymous):

1/2*1/2

OpenStudy (anonymous):

well i got -1 thts the only answer i got

OpenStudy (anonymous):

but its two answers

OpenStudy (anonymous):

why is it two answers?

OpenStudy (anonymous):

- 1 and 0.5

OpenStudy (anonymous):

then you must have done some calculation wrong. Post what you got when you substituted x = 0.5

OpenStudy (anonymous):

cant view pics, permission denied.

OpenStudy (radar):

Your reluctance to substitute the answers from using the quadratic formula is discouraging. The two answers you should of gotten were 5/2, AND 1/2. Substituting the 5/2 into your original formula would of resulted in:\[(5/2)^{2}-15/2+5/4=0\] 25/4-15/2+5/4=0 25/4-30/4+5/4=0 0=0 5/2 is a good answer. Now check out the 1/2